if you travel 52 miles in 2.5 hours whats your average speed?
Found 2 solutions by oberobic, stanbon:
Answer by oberobic(2304)

distance = rate * time
.
d = 52 miles
.
t = 2.5 hr
.
r = d/t
.
r = 52/2.5 = 20.8 mph
.
Done.

Answer by stanbon(75887)

if you travel 52 miles in 2.5 hours whats your average speed?
------
rate = distance/time
----
rate = 52/2.5 = 20 mph
==============================
Cheers,
Stan H.
===========